Problem 10-08A (Part Level Submission) Bramble Corporation sold $3,300,000, 9%, 5-year bonds on January 1, 2022. The bonds we(b) Your answer is partially correct. Try again. Prepare journal entries to record the issuance of the bonds and bond interes(c) (1) Show the balance sheet presentation for the bond issue at December 31, 2022, using the 107 selling price. BRAMBLE COR . I just need part C completed.
Answer #1
c
1
Balance Sheet (Partial)
December 31,2022
Current Liabilities
Interest payable 297000
Long term Liabilities:
Bonds payable 3300000
Add: Premium on Bonds payable 184800 3484800
2
Balance Sheet (Partial)
December 31,2022
Current Liabilities
Interest payable 297000
Long term Liabilities:
Bonds payable 3300000
Less: Discount on Bonds payable 105600 3194400
Workings:
Premium on issue 231000 =3300000*0.07
Less: Premium amortized 46200 =231000/5
Premium on Bonds payable 184800
Discount on issue 132000 =3300000*0.04
Less: Premium amortized 26400 =132000/5
Discount on Bonds payable 105600